2)     A rocket is launched straight up from Earth’s surface at 2100 m/s.

a.      By ignoring air resistance and considering the conservation of mechanical energy, determine the maximum height it reaches.

b.      What other information would be required to solve this problem in part (a) if air resistance is not ignored? How would your answer in part (a) change?

 


1
Expert's answer
2021-11-29T11:41:17-0500

(a) "mgh=mv^2/2\to h=v^2/(2g)="


"=2100^2/(2\cdot9.8)=225000\ (m)"


(b) You need to set the force of air resistance. Taking into account the force of resistance, the lifting height will be smaller.